# The Vector Grounding Problem

### v2

@@ Abstract
The remarkable performance of large language models (LLMs) on complex linguistic tasks has sparked debate about their capabilities. Unlike humans, these models learn language solely from textual data without directly interacting with the world. Yet they generate seemingly meaningful text on diverse topics. This achievement has renewed interest in the classical ‘Symbol Grounding Problem’ – the question of whether the internal representations and outputs of symbolic AI systems can possess intrinsic meaning that is not parasitic on external interpretation. Although modern LLMs compute over vectors rather than symbols, an analogous problem arises for these systems, which we call the Vector Grounding Problem. This paper has two main goals. First, we distinguish five main notions of grounding that are often conflated in the literature, and argue that only one of them, which we call referential grounding, is relevant to the Vector Grounding Problem. Second, drawing on philosophical theories of representational content, we provide two arguments for the claim that LLMs and related systems can achieve referential grounding: (1) through preference fine-tuning methods that explicitly establish world-involving functions, and (2) through pre-training alone, which in limited domains may select for internal states with world-involving content, as mechanistic interpretability research suggests. Through these pathways, LLMs can establish connections to the world sufficient for intrinsic meaning. One potentially surprising implication of our discussion is that that multimodality and embodiment are neither necessary nor sufficient to overcome the Grounding Problem.
@@ 1 Introduction
The remarkable performance of Large Language Models (LLMs) in sophisticated linguistic and cognitive tasks – once considered reliable indicators of human-like intelligence – has rekindled philosophical discussions on the nature of linguistic competence (Millière & Buckner 2024, Millière forthcoming). Despite being trained exclusively on text-based data without direct interaction with the physical world, these models exhibit the ability to generate coherent and contextually relevant paragraphs, answer questions, and solve commonsense reasoning tasks in a wide range of domains. This ability raises intriguing questions about the relationship between language, meaning, and reference in disembodied artificial systems that learn from input of raw text. [^1] For recent discussion, see Butlin (2021), Cappelen & Dever (2021), Pavlick (2022), Chalmers (2023), Mandelkern & Linzen (2024), Lederman & Mahowald (2024), Miracchi Titus (2024), Grindrod (2024), Borg (2025), Pepp (2025).
@@
In this paper, we revisit the classical ‘Symbol Grounding Problem’, originally formulated for symbolic AI systems, in light of contemporary Generative Pre-Trained Models, and especially LLMs. The core issue can be summarized as follows: if AI systems are designed to process linguistic inputs without direct interaction with the world, can their internal representations and outputs possess any meaning beyond the interpretations that we, as intelligent beings embedded in the world, project onto them? We use the term intrinsic meaning to capture this notion of interpretation-independent meaning, distinguishing it from meaning that is imposed from outside the system itself, which is thus extrinsic. [^2] This follows Harnad (1990)’s terminology. Intrinsic meaning, as we understand it here, is fully compatible with externalist views about how states and outputs come to acquire interpretation-independent meaning and reference. The intrinsic/extrinsic distinction cuts across the internalism/externalism one. We thank an anonymous referee for prompting this clarification. We argue that a similar conundrum arises for Generative Pre-Trained Models, which we dub the Vector Grounding Problem. Since these systems compute over continuous vectors rather than discrete symbols, the problem assumes a slightly different form. After outlining this new version of the classical problem, we suggest that there are strong reasons to believe that a basic form of grounding is possible in current LLMs.
@@
Specifically, we contend that the learning process of LLMs establishes functional roles for their internal states that go beyond mere linguistic prediction. Although LLMs learn from text-based data, they undergo selection pressures that can favour internal states tracking features of the world whenever extra-linguistic criteria influence the optimization objective. This can occur directly by fine-tuning LLMs on objectives that reflect extra-linguistic norms – such as factuality – or indirectly via the implicit extra-linguistic success conditions of latent tasks embedded within the training data or provided in context. When internal states that were selected to respond differentially to specific worldly conditions causally influence the generation of linguistic outputs, those outputs become grounded in the world in a way that makes them meaningful independently of human interpretation. The crucial question is thus not whether outputs can be interpreted as meaningful by humans (they obviously can), but whether the outputs themselves stand in the right kinds of relations to the world such that they would be meaningful even in the absence of external interpretation.
@@
Importantly, this paper does not address whether LLMs and related models understand language (or anything at all), engage in linguistic acts, or possess cognitive or mental properties. [^3] See Goldstein & Levinstein (2024) and Chalmers (2025) for discussions that also draw on theories of content but consider whether LLMs can be ascribed propositional attitudes, and Stoljar & Zhang (2025) for a discussion of whether LLMs can think. We do not claim that these models ‘know’ what they mean, or even what meaning is, nor that they have any conception of truth or falsity. Our sole contention is that the problem of grounding internal states and outputs in the world is solvable in principle, and in some cases arguably already in practice, for LLMs and related models. While having grounded internal states may be a crucial step toward understanding, agency, knowledge, and various other cognitive and mental capacities, it is unlikely that they are, by themselves, sufficient for those capacities. This clarification is especially important, as the current debate occasionally blurs these distinctions – a conflation we strive to avoid in this paper.
@@
The paper proceeds as follows. First, we introduce the Symbol Grounding Problem in the context of classical AI (section 2). Next, we present the Vector Grounding Problem as it applies to LLMs and related models (section 3). We then differentiate between various senses of grounding, showing that only one of them – referential grounding – is essential to the Vector Grounding Problem (section 4). Drawing on philosophical theories of representational content, we articulate a set of conditions that must be fulfilled by a system to achieve referential grounding (section 5). Finally, we argue that at least some LLMs and related multimodal models satisfy these requirements, and therefore have internal representations and outputs with intrinsic meaning (section 6).
@@ 2 The Classical Symbol Grounding Problem
Research on artificial intelligence has historically followed two competing paradigms: classical (or symbolic), and connectionist. The classical paradigm focuses on systems that represent and manipulate information using discrete symbols, based on explicit rules or procedures. Symbolic manipulation is purely syntactic; it is based on the shapes of symbols and the way in which they can be combined into complex symbolic structures (Simon & Newell 1971). In contrast, the connectionist paradigm focuses on artificial neural networks that process information in a distributed way through interconnected nodes, learning patterns from data rather than following explicit symbolic rules (Mcclelland et al. 1986).
@@
The Symbol Grounding Problem, as identified by Harnad (1990), asks how the meaning of symbols in a formal symbol system can be intrinsic to the system itself, rather than merely parasitic on the meanings in the minds of external interpreters. More precisely:
@@
Symbol Grounding Problem. How can symbol tokens, manipulated solely on the basis of their shapes according to syntactic rules, acquire meaning and refer to entities and properties in the real world – rather than merely relating other meaningless symbol tokens – without depending on the semantic interpretations provided by external meaning-bearing systems? [^4] It is common to take Harnad’s Symbol Grounding Problem as an elaboration of an earlier, highly influential criticism of AI: Searle (1980)’s Chinese Room thought experiment. Harnad (1990) himself followed this line.
Harnad (1990) provides a simple and powerful intuition pump to motivate the existence of a Symbol Grounding Problem in classical, symbolic AI: the Chinese Dictionary. Take a system, be it biological or artificial, which knows no natural language. Suppose that we want it to learn Chinese (or any other language), but the only resource made available to the system is a single-language dictionary, say a comprehensive Chinese-Chinese dictionary. Will the system be able to learn Chinese just by using the dictionary? The thought experiment invites the intuition that the system would not be able to do so. After all, the most it can do is to look up strings of symbols and see what other strings of symbols follow them. The former would be the words to be defined, the latter their definitions. However, by themselves the symbols are just meaningless shapes followed by other shapes, and it seems intuitively hopeless that the meaning of words (parts of those strings of symbols) could be learned this way. What seems to be missing is a way out from the symbolic ‘merry-go-round’: to connect those symbols in the dictionary to what they pick out in the world. But how can that be possible if the system has no access to the world, but only to further symbols?
Symbolic AI systems, according to Harnad, find themselves in this predicament. All they have access to are strings of symbols and how they follow each other. There seems to be no way for connections between such symbols and the world to be established, and thus no way for the symbols to have intrinsic meaning. Humans, who do have representations arguably acquired by means of engagement with the world, are, on the other hand, able to interpret those symbolic representations and outputs as meaningful. The meaning of the representations is hence extrinsic, depending on human interpreters. Take out the interpreter, and no meaning is to be found.
In brief, symbolic AI systems, the argument goes, manipulate symbols that are not connected to the world, that are not grounded in the world; and thereby have no intrinsic meaning. Classical AI systems, in other words, suffer from the Symbol Grounding Problem.
@@ 3 The Vector Grounding Problem
Connectionist models have come a long way since the 1980s, with deep neural networks (DNNs) emerging as powerful tools for various tasks (LeCun et al. 2015). Recent research has converged onto training deep neural networks on vast amounts of data to perform a broad range of tasks they were not explicitly designed for. Such DNNs are routinely called ‘Foundation Models’ (Bommasani et al. 2022), though we will use the more neutral label ‘Generative Pre-Trained Models’. Many among the most popular Generative Pre-Trained Models today are powered by the Transformer architecture, first introduced by Vaswani et al. (2017). Transformers use an attention mechanism to model global dependencies between sequential elements in the input, such as sentences in natural language. The general inductive biases supplied by this mechanism make Transformers suitable for many domains, including natural language processing and computer vision.
@@
Some Generative Pre-Trained Models, called Large Language Models (LLMs), are trained exclusively on text. State-of-the-art LLMs like GPT-4  and Claude have over $10^{11}$ parameters and are trained on massive corpora spanning hundred of billions of words, containing a significant subset of all written text on the internet. After training, LLMs can generate well-written and coherent paragraphs about virtually any topic. Furthermore, given explicit instructions in natural language in the input sequence, or ‘prompts’, they can perform a broad range of tasks, including: summarising long-form content, translating between multiple languages, answering complex questions, engaging in commonsense reasoning, solving simple logic and maths problems, generating code for simple programs, or explaining jokes (Srivastava et al. 2023).
@@
LLMs process input text into a sequence of tokens (i.e., full words, word-parts such as ‘-ing’, punctuation marks, and other linguistic symbols). LLMs encode each token as a vector of real-valued numbers – known as an embedding – in a high-dimensional vector space, and estimate the probabilities of all possible next tokens in the sequence based on the previous embeddings. The resulting prediction of the next token in the sequence is then compared to the actual next token, and the model’s parameters are subsequently updated to make the correct token more probable in that context, and other tokens less probable. While LLMs do not treat words as the main units of processing, the combination of tokens can form an indefinite number of possible words, and we can view the vectors for such composing tokens as jointly forming a word embedding. This training process (confusingly known as pre-training) encourages the model to produce context-sensitive embeddings, with semantically similar words having vectors that are close together in the high-dimensional space.
@@
There is a deep connection between the way in which LLMs learn embeddings from linguistic corpora, and the old idea from structural linguistics according to which word meaning can be inferred from the contexts in which they appear, known as the ‘distributional hypothesis’ (Harris 1954, Firth 1957). In LLMs, semantic similarity is determined based on distribution across contexts: words that appear in similar contexts are modelled by nearby vectors. The dimensions of the vectors are tuned by the model to best predict surrounding words, and capture distributional patterns. Although individual dimensions are not typically interpretable, the positions of vectors in the space are a result of statistical patterns in the linguistic data, and reflect semantic similarities in light of how the words are used in context. [^5] Vision-Language Models (VLM) that can process images in addition to text work in a similar way: they encode text and images as vectors in a single high-dimensional space. We will discuss whether these models have specific implications for the grounding problem in subsection 7.3. In this way, embeddings in LLMs are not arbitrary or conventional – as the symbols in classical AI systems typically are – since they are shaped by statistical patterns in language.
@@
However, LLMs appear to face a similar predicament to that of traditional symbolic AI systems when it comes to grounding and intrinsic meaning. Their reliance on continuous vectors, rather than discrete symbols, seems to make no difference to escaping the merry-go-round of representations, and being connected to the external world – LLMs are trapped in a carousel of numbers, rather than symbols. Much like their classical counterparts, LLMs manipulate vectors that do not seem grounded in the world and thus lack intrinsic meaning. Consequently, they grapple with an analogous difficulty, which we call the Vector Grounding Problem.
Mirroring Harnad’s 1990 Symbol Grounding Problem, the Vector Grounding Problem asks how the vector embeddings inside Generative Pre-Trained Models can have intrinsic meanings, rather than merely being parasitic on the meanings in the minds of external interpreters. More precisely:
@@
Vector Grounding Problem. How can vector embeddings, manipulated solely based on their numerical values according to algebraic operations, acquire meaning and refer to entities and properties in the real world – rather than merely relating other meaningless vectors – without depending on the semantic interpretations provided by external meaning-bearing systems?
@@
Both Grounding Problems, as stated here, focus on internal states, such as symbols and vector embeddings. If the Grounding Problems are solved, and such internal states come to have intrinsic meaning, we get straightforward reasons to think that the systems in question produce outputs that are intrinsically meaningful. For the meaning of the outputs would thus derive from the meaning of the system’s internal representations, rather than from the interpretations given by human users. [^6] In such a case, human users could misinterpret the system’s output if their interpretation does not conform to what the system actually represented – in a way akin to how humans can misinterpret each other.
@@
It is worth noting that the Grounding Problems are particularly relevant to systems that generate novel outputs. This sets them apart from systems that merely reproduce outputs generated by humans, such as paper copiers. When a human-generated text is copied, there is no question about its intrinsic meaning: it was produced by a system with meaningful internal representations to convey a specific meaning to the reader. In contrast, generative systems like LLMs produce outputs that are not, for the most part, mere copies of human-generated text; they generate entirely novel sentences whose meanings cannot be equated to what any specific human has ever thought or said. Unlike with paper copiers, the question of whether LLM outputs are meaningful cannot be straightforwardly explained by appealing to a human author. [^7] We thank an anonymous referee for pressing us on this point.
@@
Bender & Koller (2020) propose a thought-provoking thought experiment, the “Octopus Test”, to pump intuitions about the inescapability of the Vector Grounding Problem for LLMs:
@@
The Octopus Test. Two English-speaking castaways find themselves stranded on neighbouring islands, separated by treacherous waters. Fortuitously, they discover telegraphs left by previous inhabitants, connected via an underwater cable, which enables them to communicate by exchanging telegraphic messages. Unbeknownst to them, a superintelligent octopus inhabits these waters and taps into the underwater cable, intercepting their messages. Though the octopus lacks any knowledge of English, its superintelligence allows it to detect statistical patterns in the telegraphic messages and to form an accurate representation of the statistical relationships between various telegraphic signals. The octopus decides to sever the underwater cable and position itself at the ends of the two resulting cable segments, receiving and replying to telegraphic signals from both castaways based on the statistical patterns it has identified. Whether or not the castaways notice this change, the messages sent by the octopus intuitively seem to hold no intrinsic meaning. After all, the octopus simply adheres to the statistical patterns it has learned from listening in on the previous exchanges between the humans, without any understanding of the human interpretation of the signals, such as ‘coconut’ or ‘tree’. Furthermore, the octopus likely does not comprehend that the signals possess meaning or serve a communicative function. [^8] Bender & Koller (2020) emphasise the fact that the octopus lacks communicative intent when generating telegraphic signals. Since our focus here is the narrower, more basic question of vector grounding, rather than the more complex capacities related to language comprehension and communication, we will set those considerations aside.
@@
As should be clear, the octopus in this thought experiment is operating roughly as an extremely powerful LLM would. Regardless of how the octopus represents the signals and their statistical relations, be it through vectors or other means, it only has access to the signals and their statistical relations – without any connection to the entities and properties in the real world that the words encoded in those signals represent. Consequently, even if the octopus can produce outputs that appear sensible to humans in all situations, these outputs lack intrinsic meaning. [^9] Our version of the Octopus Test goes beyond Bender & Koller (2020) by assuming that it is possible the humans would never notice they were no longer communicating with each other. In contrast, Bender & Koller (2020) contend that the octopus would fail to produce appropriate outputs to questions involving concrete objects it has never perceived, such as ‘how do I build a coconut catapult?’. This claim seems to be challenged by the often plausible replies that current LLMs provide to this sort of question. We should expect the superintelligent octopus to generate even more reliably plausible responses. The Octopus Test implies that LLMs generate outputs as devoid of intrinsic meaning as those produced by the superintelligent octopus. [^10] Arguably, LLMs’ outputs would be even more devoid of intrinsic meaning than the octopus’, given that the latter is a biological organism with cognitive states, motivation, superintelligence, and other features that LLMs lack. Similar considerations apply to the comparison between LLMs and parrots (Bender et al. 2021). In the case of these animals, even if their outputs carry some meaning, they typically do not align with the meaning of the language they produce. It is humans, whether castaways or otherwise, who interpret these outputs as meaningful because they resemble meaningful words and sentences. LLMs have no access to coconuts, trees, or anything else beyond language. The vectors they receive and manipulate during internal processing are merely meaningless arrays of numbers, ungrounded in anything outside statistical patterns in language, just like the outputs they generate. In other words, the Octopus Test invites the intuition that all meaning in the outputs produced by LLMs is extrinsic, i.e., it lies in the eye of the beholder, their human interpreters. This would be so, we might add, because LLMs are susceptible to the Vector Grounding Problem.
@@
Throughout the remainder of this paper, we will challenge this claim and counter the intuition primed by the Octopus Test. We aim to show that, contrary to first appearances, LLMs indeed possess the features required to generate meaningful outputs based on meaningful internal representations.
Before developing this argument further, it is essential to clarify the target notion of grounding.

### v3

@@ Abstract
Large language models (LLMs) produce seemingly meaningful outputs, yet they are trained on text alone without direct interaction with the world. This leads to a modern variant of the classical symbol grounding problem in AI: can LLMs’ internal states and outputs be about extra-linguistic reality, independently of the meaning human interpreters project onto them? We argue that they can. We first distinguish referential grounding—the connection between a representation and its worldly referent—from other forms of grounding and argue it is the only kind essential to solving the problem. We contend that referential grounding is achieved when a system’s internal states satisfy two conditions derived from teleosemantic theories of representation: (1) they stand in appropriate causal-informational relations to the world, and (2) they have a history of selection that has endowed them with the function of carrying this information. We argue that LLMs can meet both conditions, even without multimodality or embodiment.
@@ 1 Introduction
The remarkable performance of Large Language Models (LLMs) in sophisticated linguistic and cognitive tasks—once considered reliable indicators of human-like intelligence—has rekindled philosophical discussions on the nature of linguistic competence. [^1] See Milliere2024 for a general introduction to language models, and Milliereforthcoming for a discussion of their relevance to debates about linguistic competence Despite being trained exclusively on text data without direct interaction with the physical world, these models exhibit the ability to generate coherent and contextually relevant paragraphs, answer questions, and solve commonsense reasoning tasks in a wide range of domains. This ability raises intriguing questions about the relationship between language, meaning, and reference. [^2] For recent discussion, see Butlin2021, cappelenMakingAIIntelligible2021, Pavlick_2022, chalmersDoesThoughtRequire2023, mandelkernLanguageModelsWords2024, Lederman2024, miracchititusDoesChatGPTHave2024, grindrodLargeLanguageModels2024, borgLLMsTuringTests, peppReferenceIntentionsLarge.
@@
In this paper, we revisit the classical ‘Symbol Grounding Problem’ (Harnad1990), originally formulated for symbolic AI systems, in light of contemporary LLMs. The core issue can be summarized as follows: if AI systems process linguistic inputs without direct interaction with the world, can their internal representations and outputs possess any meaning beyond the interpretations that we, as intelligent beings embedded in the world, project onto them? We use the term intrinsic meaning to capture this notion of interpretation-independent meaning, distinguishing it from meaning that is imposed from outside the system itself, which is thus extrinsic. [^3] This follows Harnad1990’s terminology. Intrinsic meaning, as we understand it here, is fully compatible with externalist views about how states and outputs come to acquire interpretation-independent meaning and reference. The intrinsic/extrinsic distinction cuts across the internalism/externalism one. We argue that a similar conundrum arises for LLMs, which we call the Vector Grounding Problem. Since these systems compute over continuous vectors rather than discrete symbols, the problem takes a somewhat different form. After outlining this new version of the classical problem, we suggest that there are strong reasons to believe that a basic form of grounding is possible in current LLMs.
@@
Specifically, we contend that the learning process of LLMs endows some of their internal states with functions that go beyond mere linguistic prediction. Although LLMs learn from text data, they undergo selection processes that favour internal states tracking features of the world whenever extra-linguistic criteria influence the training objective. This can occur directly by fine-tuning LLMs on objectives that reflect extra-linguistic norms—such as factuality—or indirectly via the implicit extra-linguistic success conditions of latent tasks embedded within the training data or provided in context. When internal states selected to carry information about specific worldly conditions causally influence the generation of linguistic outputs, those outputs become grounded in the world in a way that makes them meaningful independently of human interpretation. The crucial question is thus not whether outputs can be interpreted as meaningful by humans (they obviously can), but whether the outputs themselves stand in the right kinds of relations to the world such that they would be meaningful even in the absence of external interpretation.
@@
Importantly, this paper does not address whether LLMs and related models understand language (or anything at all), engage in linguistic acts, or possess cognitive or mental properties. [^4] See goldsteinDoesChatGPTHave2024 and chalmersPropositionalInterpretabilityArtificial2025 for discussions that also draw on theories of content but consider whether LLMs can be ascribed propositional attitudes, and stoljarWhyChatGPTDoesnt2025 for a discussion of whether LLMs can think. We do not claim that these models ‘know’ what they mean, or even what meaning is, nor that they have any conception of truth or falsity. Our sole contention is that the problem of grounding internal states and outputs in the world is solvable in principle, and in some cases arguably already in practice, for LLMs and related models. While having grounded internal states may be a crucial step toward understanding, agency, knowledge, and various other cognitive and mental capacities, it is unlikely that such states are, by themselves, sufficient for those capacities. This clarification is especially important, as the current debate occasionally blurs these distinctions—a conflation we strive to avoid in this paper.
@@
We proceed as follows. First, we introduce the Symbol Grounding Problem in the context of classical AI (section 2). Next, we present the Vector Grounding Problem as it applies to LLMs and related models (section 3). We then differentiate between various notions of grounding, showing that only one of them—referential grounding—is essential to the Vector Grounding Problem (section 4). Drawing on philosophical theories of representational content, we articulate a set of conditions that a system must satisfy to achieve referential grounding (section 5). We argue that at least some LLMs and related multimodal models satisfy these requirements, and therefore have internal representations and outputs with intrinsic meaning (section 6). Finally, we discuss some surprising implications of our argument (section 7).
@@ 2 The Classical Symbol Grounding Problem
Research in artificial intelligence has historically been shaped by two distinct paradigms: the classical (or symbolic) approach and the connectionist approach. The classical paradigm focuses on systems that represent and manipulate information through discrete symbols according to explicit rules. This symbolic manipulation is purely syntactic; it depends solely on the formal properties of symbols and the rules that govern their combination into complex structures (simonHumanProblemSolving1971). By contrast, the connectionist paradigm uses artificial neural networks that process information through massively parallel and distributed computation across interconnected nodes. Rather than following predetermined symbolic rules, these systems induce patterns directly from data through learning algorithms (mcclellandParallelDistributedProcessing1986).
@@
The Symbol Grounding Problem, as identified by Harnad1990, asks how the meaning of symbols in a formal symbol system can be intrinsic to the system itself, rather than merely parasitic on the meanings in the minds of external interpreters. More precisely:
@@
Symbol Grounding Problem. How can symbol tokens, manipulated solely on the basis of their shapes according to syntactic rules, acquire meaning and refer to entities and properties in the real world—rather than merely relating to other meaningless symbol tokens—without depending on the semantic interpretations provided by external meaning-bearing systems? [^5] It is common to take Harnad’s Symbol Grounding Problem as an elaboration of an earlier, highly influential criticism of AI: Searle1980’s Chinese Room thought experiment. Harnad1990 himself followed this line.
Harnad1990 provides a compelling intuition pump to motivate the Symbol Grounding Problem in classical symbolic AI. Consider a system—biological or artificial—with no prior knowledge of any natural language. Suppose we attempt to teach this system Chinese (or any other language), but provide it with only a monolingual resource: a comprehensive Chinese-Chinese dictionary. Can the system learn Chinese from this dictionary alone? The thought experiment invites a negative answer. The system can look up strings of symbols and find other strings that purportedly define them—words paired with their definitions. However, these symbols remain mere meaningless shapes followed by other shapes. Without any connection to the external world, it seems impossible for the system to grasp the actual meaning of these words. The system is trapped in a symbolic ‘merry-go-round’, endlessly cycling through symbols that refer only to other symbols, never breaking free to establish connections with real-world referents. Without access to anything beyond symbols, how can the system ever ground these symbols in the real world?
According to Harnad, symbolic AI systems find themselves in this predicament. Limited to manipulating strings of symbols and their syntactic relationships, these systems lack any mechanism for establishing connections between symbols and their real-world referents. Consequently, the symbols possess no intrinsic meaning. While humans—who do have representations acquired through direct engagement with the world—can interpret these symbolic outputs as meaningful, this meaning remains extrinsic, entirely dependent on human interpretation. Remove the human interpreter, and the symbols become meaningless shapes. In essence, the argument shows that symbolic AI systems manipulate ungrounded symbols: symbols disconnected from the world they purport to represent. These systems thus lack intrinsic semantic content. Classical AI systems, in other words, are fundamentally limited by the Symbol Grounding Problem.
@@ 3 The Vector Grounding Problem
Connectionist models have come a long way since the 1980s, with deep neural networks (DNNs) emerging as powerful tools for various tasks (lecunDeepLearning2015). Recent research has converged on training deep neural networks on vast amounts of data to perform a broad range of tasks they were not explicitly designed for. Such DNNs are often called ‘foundation models’ (bommasaniOpportunitiesRisksFoundation2022), though we will use the more descriptive label ‘generative pre-trained models’. Most of these models are built with a neural network architecture called the ‘Transformer’, introduced by Vaswani2017, which uses a mechanism called ‘attention’ to capture global dependencies between sequential elements. This architecture has proved particularly effective for linguistic and visual data.
@@
LLMs are generative pre-trained models trained on text data. State-of-the-art LLMs have over $10^{11}$ parameters and are trained on massive corpora spanning trillions of words, effectively ingesting a substantial fraction of all text available on the internet. After training, LLMs can generate well-written and coherent paragraphs about virtually any topic. Furthermore, given explicit instructions in natural language in the input sequence, or ‘prompts’, they can perform a broad range of tasks, including: summarising long-form content, translating between multiple languages, answering complex questions, engaging in commonsense reasoning, solving logic and maths problems, generating code, or explaining jokes (srivastava2023beyond).
@@
LLMs process text into a sequence of tokens: units that may map onto whole words or word parts. Each token is encoded as a high-dimensional vector called an embedding. The model then estimates a probability distribution over possible next tokens based on the sequence of token embeddings. During training, next-token predictions are compared against the actual next token in the training data, and model parameters are adjusted to increase the probability of correct tokens while suppressing alternatives. Though LLMs do not treat words as the main units of processing, the combination of tokens can form an indefinite number of possible words, and we can view the vectors for such composing tokens as jointly forming a word embedding. This training process (confusingly known as pre-training) encourages the model to produce context-sensitive embeddings, with semantically similar words having vectors that are close together in the high-dimensional space.
@@
There is a deep connection between the way in which LLMs learn embeddings from linguistic corpora, and the old idea from structural linguistics according to which word meaning can be inferred from the contexts in which they appear, known as the distributional hypothesis (harrisDistributionalStructure1954, firthPapersLinguistics193419511957). In LLMs, semantic similarity is determined based on distribution across contexts: words that appear in similar contexts are mapped to nearby vectors. Although individual dimensions of the vector space are not typically interpretable, the geometric relationships between vectors encode statistical regularities that reflect latent semantic structure. Consequently, embeddings are not purely conventional, unlike symbols in classical AI systems, as their relational structure is determined by statistical patterns in linguistic data. [^6] Vision-Language Models (VLMs) extend this approach by encoding both textual and visual inputs within a unified vector space. We examine the implications of such multimodal architectures for the grounding problem in subsection 7.3.
@@
Nevertheless, LLMs seem to face a challenge analogous to that of traditional symbolic AI systems regarding grounding and intrinsic meaning. Their use of continuous vectors rather than discrete symbols does not help them escape the ‘merry-go-round’ of representations or connect to the external world—LLMs seem trapped in a carousel of numbers instead of symbols. Like their classical counterparts, LLMs manipulate vectors that appear to lack grounding in the world and thus have no intrinsic meaning. Consequently, they encounter an analogous predicament, which we call the Vector Grounding Problem:
@@
Vector Grounding Problem. How can vector embeddings, manipulated solely on the basis of their numerical values according to algebraic operations, acquire meaning and refer to entities and properties in the real world—rather than merely relating to other meaningless vectors—without depending on the semantic interpretations provided by external meaning-bearing systems?
@@
Both versions of the Grounding Problem, as stated here, focus on the semantic content of internal states such as symbols and vector embeddings. If either problem is solved, so that their internal states acquire intrinsic meaning, we would have compelling grounds for attributing genuine semantic content to the system’s outputs: the meaning of outputs would derive from the system’s internal representations, rather than from post-hoc interpretations by human users. [^7] This would create the possibility of genuine misinterpretation: human users could misconstrue the system’s output by failing to grasp what the system actually represented, analogous to how humans routinely misinterpret each other’s statements.
@@
The Grounding Problems are particularly relevant to systems that generate novel outputs. This sets them apart from systems that merely reproduce outputs generated by humans, such as paper copiers. When a human-generated text is copied, there is no question about its intrinsic meaning: it was produced by a system with meaningful internal representations to convey a specific meaning to the reader. In contrast, generative systems like LLMs produce outputs that are not, for the most part, mere copies of human-generated text; they generate novel sentences whose meanings cannot be equated to what any specific human has ever thought or said. Unlike with paper copiers, the question of whether LLM outputs are meaningful cannot be straightforwardly explained by appealing to a human author. [^8] We will return to the question of whether the representations of LLMs only acquire extra-linguistic content through artefactual functions that depend on the intentions of their human designers in section 6.2.1.
@@
Bender2020 propose a thought experiment, the ‘Octopus Test’, to pump intuitions about the inescapability of the Vector Grounding Problem for LLMs:
@@
Octopus Test. Two English-speaking castaways find themselves stranded on neighbouring islands, separated by treacherous waters. Fortuitously, they discover telegraphs left by previous inhabitants, connected via an underwater cable, enabling them to communicate through telegraphic messages. Unbeknownst to them, a superintelligent octopus inhabiting these waters taps into the cable, intercepting their messages. Though the octopus has no knowledge of English, its superintelligence enables it to detect statistical patterns in the telegraphic exchanges and construct an accurate model of the statistical relationships between various signals. The octopus then severs the cable and positions itself at both ends, intercepting messages from each castaway and generating responses based solely on the statistical patterns it has learned. Whether or not the castaways detect this substitution, the octopus’s messages intuitively seem devoid of intrinsic meaning. After all, the octopus merely reproduces statistical patterns gleaned from eavesdropping on human exchanges, without understanding what the signals denote. Moreover, the octopus likely fails to grasp that these signals carry meaning or serve any communicative function. [^9] Bender2020 emphasize that the octopus lacks communicative intent when generating telegraphic signals. Since our focus is on the narrower question of vector grounding rather than the broader capacities underlying language comprehension and communication, we set aside these considerations.
@@
The octopus in this thought experiment operates analogously to a powerful LLM. Regardless of the format of its internal states—whether vectors or otherwise—the octopus accesses only linguistic signals and their statistical relationships, lacking any connection to the real-world entities and properties these signals purportedly represent. Consequently, even if the octopus produces outputs that appear perfectly sensible to human observers, these outputs lack intrinsic meaning. The Octopus Test suggests that LLMs generate outputs as devoid of intrinsic meaning as those produced by the superintelligent octopus. [^10] One might argue that LLMs’ outputs are even more devoid of intrinsic meaning than the octopus’s, given that the latter is a biological organism endowed with cognitive states, motivations, superintelligence, and other features that LLMs lack. Similar considerations apply when comparing LLMs to parrots (Bender2021). In the case of these animals, even if their outputs carry some meaning, they typically do not align with the meaning of the language they produce. Humans (castaways or contemporary users) project meaning onto these outputs because they superficially resemble meaningful words and sentences. LLMs have no access to coconuts, trees, or any extra-linguistic reality. The vectors they process are merely arrays of numbers, grounded only in statistical patterns within language, not in the world these patterns purportedly describe. In other words, the Octopus Test invites the conclusion that all meaning in the outputs produced by LLMs is extrinsic: it lies in the eye of the beholder. This would be because LLMs are susceptible to the Vector Grounding Problem.
@@
Throughout the remainder of this paper, we challenge this claim and the intuitions primed by the Octopus Test. We argue that, contrary to initial appearances, LLMs possess the necessary features to generate meaningful outputs based on internal representations with extra-linguistic content. Before developing this argument, we must first clarify our target notion of grounding.